### v1

@@
Abstract
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can raise electricity demand beyond the safe limits of electrical infrastructure, increasing the risk of blackouts or requiring grid reinforcement that can be slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to three times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles could require up to 312 GW of distribution grid reinforcement nationally, at a cost of \$183 to \$415 billion, or \$1,500 to \$3,400 per household (95% confidence intervals). However, demand-side management can mitigate demand peaks, reducing grid reinforcement costs by up to 92%.
keywords:
@@ Abstract
Fossil-fueled driving, space heating, and water heating contribute about one-third of United States greenhouse gas emissions [1]. Replacing fossil-fueled vehicles and heaters with electric alternatives could significantly reduce these emissions today [2], and essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 [3]. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates [4, 5, 6]. State and local governments have enacted further incentives, including some bans on fossil-fueled vehicles or heaters [7]. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating [8].
@@
While electrification of fossil-fueled activities could deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning [9]. Electrifying space heating could heighten annual demand peaks and shift them from summer to winter [10, 11, 12, 13, 14]. Electric vehicle adoption could further exacerbate demand peaks [15, 16, 17]. Heightened demand peaks from electrification could require replacing power lines and transformers, which must be sized to accommodate worst-case power flows. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years [18]. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices [19, 20], which could weaken economic incentives for further electrification.
@@
Existing research on nationwide grid impacts of electrification focuses mainly on high-voltage transmission systems. The National Renewable Energy Laboratory’s Electrification Futures Study [21, 10, 22] investigates how electrification might alter load profiles and transmission infrastructure needs. In their high-electrification 2050 scenario, peak electricity demand rises 38% relative to a reference scenario. Wilson et al. [2] use the ResStock tool [23] to investigate cost and emission impacts under various heat pump adoption scenarios. While [2] does not explicitly consider grid impacts, Wilson et al. stress the need to understand how evolving load profiles should shape grid infrastructure investments.
@@
In parallel with nationwide transmission-level research, several studies investigate distribution-level grid impacts in California. Notably, Elmallah et al. [24], use the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimate that distribution grid reinforcement to accommodate new load from electric vehicles and heat pumps could cost \$1 to \$10 billion, or about \$200 to \$2,000 per customer. Li et al. [25] use the same dataset to study electric vehicle adoption statewide, estimating that 67% of distribution feeders will need reinforcement, at a total cost of \$6 to \$20 billion. Several other state-specific studies have investigated the impacts of electrification on emissions [14, 26, 27].
As a complement to nationwide transmission-level studies and California-specific distribution-level studies, this paper investigates the impacts of electrification of space heating, water heating, and personal vehicles on distribution grids nationwide. We use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox [28] and make all data publicly available to facilitate reproduction or extension of the work. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather [29], miscellaneous electrical loads [30], indoor temperature setpoints [31], and domestic hot water use [32]. EDGIE implements empirically validated, multi-physics modeling at hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle properties to county-specific data. We find that almost all United States counties could require grid reinforcement in a 100% electrification scenario. In colder climates, peak demand rises by 180% on average relative to today’s peak demand, and up to 300% in the coldest counties. Distribution grid reinforcement to accommodate 100% electrification nationwide could cost \$183 to \$415 billion. However, demand-side management strategies, such as improving equipment efficiencies and coordinating device operation, could reduce nationwide grid reinforcement costs by up to 92%.
@@
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types (detached single-family vs. attached multifamily), and levels of insulation and air sealing all vary both across and within counties. We simulate peak heating and cooling weeks under two scenarios: (1) A business-as-usual scenario that reflects today’s building stock, appliances, and vehicles, and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a resistance water heater, an air-to-air heat pump for space heating and cooling, and backup resistance heat. We compute design heating and cooling loads for each home under EnergyStar heating and cooling design conditions [33]. Each heat pump is sized to meet the larger of its design heating load and design cooling load, capped at the largest residential air-to-air heat pump size currently available. In the two northernmost climate zones, we simulate adoption of heat pumps that meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications [34]. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. Section Methods and the Supplementary Information document provide more details on models and data sources.
@@
The modeling approach outlined above enables analysis of the sizes and causes of likely grid impacts from electrification. Fig. 2 shows the aggregate power used by 1,000 households over a peak cooling week in Miami, Florida (left plot; a hot, humid climate) and a peak heating week in Minneapolis, Minnesota (right; cold). In these plots, the shaded areas from bottom to top represent miscellaneous electrical loads (blue; lights, computers, dishwashers, etc.), water heaters (orange), electric vehicles (yellow), and heat pumps with backup resistance heat (purple). Dashed red lines show today’s annual demand peaks. In Miami, electrification increases peak demand only modestly. Electric vehicle adoption increases electricity demand, but the new air-to-air heat pumps – which can both heat and cool – offset this increase substantially by improving air conditioning efficiency. In Minneapolis, by contrast, peak demand more than doubles due to space heating.
Figure 3: Today (top left), per-household demand peaks are typically higher in hotter areas. In an all-electric future (top right), per-household peaks increase and shift to colder areas. Per-household grid reinforcement requirements (bottom) are highest in the Rocky Mountains, northern Midwest, and Northeast.
@@
The lower heat map in Fig. 3 shows the estimated grid reinforcement capacity, normalized by the number of homes, required to accommodate 100% electrification in each county. We estimate the grid reinforcement capacity requirement by the difference between the all-electric and business-as-usual grid capacities. As expected, northern regions require more grid reinforcement due to higher space heating demand. Grid reinforcement requirements reach a maximum of 21 kW per household in the northern Midwest. Many southern areas with mild winters require little or no grid reinforcement. In some cases, the lower heat map in Fig. 3 shows significant variation in grid reinforcement requirements between adjacent counties. This variation can be explained by differences in housing types and sizes, levels of insulation and air sealing, driving patterns, weather extremes, or other location-specific data. These geographic differences suggest a need for tailored infrastructure planning, as approaches that work in one county may not necessarily work in neighboring counties.
@@
To assess the economic costs of reinforcing distribution grids to accommodate electrification, we multiply the aggregate grid reinforcement requirement for each county (in units of kW) by a \$/kW price. Informed by several recent studies, we choose this price to reflect a typical distribution grid upgrade. In reality, grid reinforcement prices vary significantly across the United States due to variation in grid topology and capacity, prevalence of overhead vs. underground power lines, costs of labor and equipment, and other factors. Due to a lack of reliable location-specific data on grid reinforcement prices, however, we use the same price distribution for each county.
We draw on four studies [24, 35, 36, 37] to determine a plausible range of prices for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. [24] estimate prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects [^1] We inflation-adjusted all price data to 2024 dollars.. Larson et al. [35] use Energy Information Administration data [38] to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Fares et al. [36] estimate an average price of 69 \$/kW based on data from the Federal Energy Regulatory Commission’s Form 1, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. In estimating this price, Fares et al. delineate distribution grid reinforcements from generation, transmission, and administrative costs. Rauschkolb et al. [37] use the same Federal Energy Regulatory Commission dataset to develop a linear model that accounts for both one-time capital costs of expanding grid capacity, and for increases in the annual costs of operating and maintaining a larger grid.
Here, we modify the approach of Rauschkolb et al. to estimate a representative distribution grid reinforcement price. We assume that grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. As in [37], the modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Supplementary Information document contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
The upper row of heat maps in Fig. 4 shows the estimated total distribution grid reinforcement cost for each of the lower 48 United States under low (left), medium (middle), and high (right) price scenarios. We compute these estimates by multiplying the grid reinforcement requirement, aggregated over all counties in the state, by prices of 587, 960, and 1,331 \$/kW, corresponding to the 2.5th, 50th, and 97.5th price percentiles. Unlike the per-household grid reinforcement requirement estimates in Fig. 3, the aggregate cost estimates in Fig. 4 reflect population density as well as load growth intensity. States with both high populations and high per-household grid reinforcement requirements have the highest total costs. For example, the 7.6 million households in New York, a densely populated state with cold winters, see a total cost of about \$30 billion, or about \$4,000 per household. Per-household grid reinforcement costs, shown in the bottom row of heat maps in Fig. 4, reach a maximum of \$16,000 per household under the 97.5th price percentile in the cold, sparsely populated state of North Dakota.
Aggregated over the lower 48 United States, grid reinforcement to accommodate electrification could cost \$183 to \$415 billion, or \$1,500 to \$3,400 per household. For context, United States households spend about \$233 billion (\$1,900 per household) per year on residential energy bills [39]. The Inflation Reduction Act of 2022 allocates about \$370 billion over ten years to spending on energy and climate [40]. The United States annual military budget was about \$816 billion in 2023 [41].
@@
The estimated distribution grid reinforcement requirements and costs in the previous sections come from an electrification scenario that extrapolates forward today’s housing stock, vehicle fleet, equipment configurations and efficiencies, and user behavior. Unmodified, these electrification choices could require \$183 to \$415 billion in distribution grid reinforcement costs. However, electrification with strategic demand-side management – referred to here as ‘smart electrification’ – could reduce these costs substantially.
To assess the cost reduction potential of smart electrification, we simulate four demand-side management strategies. First, the resistance water heaters simulated above could shift to heat-pump water heaters, which use about half the electricity per unit of heat output. Second, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached single-family housing to less energy-intensive attached multifamily housing, or due to shifts to smaller housing. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. We simulate device coordination through convex optimization with perfect forecasts in one representative location per state, as discussed in the Supplementary Information document. Fourth, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat.
Figure 5: Improving equipment efficiencies (top left, bottom right), reducing thermal demand (top right), or coordinating device operation (bottom left) can reduce the grid impacts of electrification.
@@
Fig. 6 shows nationwide grid reinforcement cost reductions from the four individual strategies and from combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$300 billion without demand-side management. Individually (top set of bars), the four strategies reduce costs by 6%, 22%, 52%, and 55%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication – achieves similar cost reductions to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of heat-pump water heaters (6% individually), less thermal demand from buildings (22%), and ground-source heat pumps (55%) reduces costs by 68%. The combination of all four demand-side management strategies reduces nationwide grid reinforcement costs by 92%, to \$25 billion.

### v2

@@
Distribution Grids May Be a Barrier to Residential Electrification
Priyadarshan Mechanical Engineering, Purdue University, USA Constance Crozier Industrial and Systems Engineering, Georgia Institute of Technology, USA Kyri Baker Civil, Environmental, and Architectural Engineering, University of Colorado Boulder, USA Renewable and Sustainable Energy Institute, University of Colorado Boulder, USA Kevin J. Kircher Mechanical Engineering, Purdue University, USA Electrical and Computer Engineering, Purdue University, USA Lead contact Correspondence: kircher@purdue.edu
SUMMARY
Replacing fossil-fueled appliances and vehicles with electric alternatives can reduce greenhouse gas emissions and air pollution in many settings. However, residential electrification can raise electricity demand beyond the safe limits of electrical infrastructure, increasing the risk of blackouts or requiring grid reinforcement that can be slow and expensive. Here, we estimate the physical and economic impacts on distribution grids of electrifying all housing and personal vehicles in each county of the lower 48 United States. We find that space heating is the main driver of grid impacts, with the coldest regions seeing demand peaks up to five times higher than today’s peaks. Accommodating electrification of all housing and personal vehicles could require up to 544 GW of distribution grid reinforcement nationally, at a cost of \$320 to \$720 billion, or \$2,600 to \$5,800 per household (95% confidence intervals). However, demand-side management can mitigate demand peaks, reducing grid reinforcement costs by up to 77%.
KEYWORDS
@@ INTRODUCTION
Fossil-fueled driving, space heating, and water heating contribute about one-third of United States greenhouse gas emissions 1. Replacing fossil-fueled vehicles and heaters with electric alternatives could significantly reduce these emissions today 2, and essentially eliminate them if the United States achieves its goal of 100% carbon-free electricity by 2035 3. For these reasons, the Inflation Reduction Act of 2022 incentivizes adoption of electric vehicles and heat pumps (efficient electric heating and cooling machines) through tax credits and point-of-sale rebates 4, 5, 6. State and local governments have enacted further incentives, including some bans on fossil-fueled vehicles or heaters 7. Due in part to these policies, adoption of electric vehicles and heat pumps in the United States is accelerating 8.
@@
While electrification of fossil-fueled activities could deeply reduce emissions, it could also have serious impacts on the power grid. Today, electricity demand in most United States power systems peaks in the hottest hours of the year, driven mainly by air conditioning 9. Electrifying space heating could heighten annual demand peaks and shift them from summer to winter 10, 11, 12, 13, 14. Electric vehicle adoption could further exacerbate demand peaks 15, 16, 17. Heightened demand peaks from electrification could require replacing power lines and transformers, which must be sized to accommodate worst-case power flows. Transformers in particular are already in short supply, with United States utilities experiencing lead times of up to two years and prices that have increased by 400 to 900% in the last three years 18. Utilities pass grid reinforcement costs on to ratepayers in the form of increased electricity prices 19, 20, which could weaken economic incentives for further electrification.
@@
Existing research on nationwide grid impacts of electrification focuses mainly on high-voltage transmission systems. The National Renewable Energy Laboratory’s Electrification Futures Study 21, 10, 22 investigates how electrification might alter load profiles and transmission infrastructure needs. In their high-electrification 2050 scenario, peak electricity demand rises 38% relative to a reference scenario. Wilson et al. 2 use the ResStock tool 23 to investigate cost and emission impacts under various heat pump adoption scenarios. While Wilson et al. 2 do not explicitly consider grid impacts, they stress the need for future work to understand how evolving load profiles should shape infrastructure investments.
@@
In parallel with nationwide high-voltage transmission grid research, several studies investigate medium-voltage distribution grid impacts in California specifically. Notably, Elmallah et al. 24 use the Distribution Deferral Opportunity Report, which the California Public Utilities Commission requires utilities to file each year, to analyze electrification impacts within the Pacific Gas & Electric service territory. Elmallah et al. estimate that distribution grid reinforcement to accommodate new load from electrification, including up to 100% adoption of electric vehicles and up to 50% adoption of heat pumps and water heaters, could cost \$1 to \$10 billion, or about \$200 to \$2,000 per customer. Li et al. 25 use the same dataset to study adoption of electric vehicles (but not heat pumps or water heaters) statewide, estimating that 67% of distribution feeders will need reinforcement, at a total cost of \$6 to \$20 billion or about \$300 to \$900 per customer. Several other state-specific studies have investigated the impacts of electrification on greenhouse gas emissions without considering impacts on power grids 14, 26, 27.
@@
As a complement to nationwide transmission-level studies and California-specific distribution-level studies, this paper investigates the impacts of electrification on distribution grids nationwide. In addition to expanding the scope of distribution-level studies from the 58 counties in California to 3,100 counties in the lower 48 United States, this paper investigates strategies to mitigate grid impacts through demand-side management of space heating, water heating, and personal vehicles. These are the three main drivers of grid impacts from electrification, and we are aware of no distribution-level study that has considered demand-side management strategies for all of them. We use the open-source EDGIE (Emulating the Distribution Grid Impacts of Electrification) toolbox 28 and make all data publicly available to facilitate reproduction or extension of the work. As illustrated in Fig. 1, EDGIE models heat pumps, electric vehicles, water heaters, and building thermal dynamics. It uses real data on weather 29, indoor temperature setpoints 30, domestic hot water use 31, and miscellaneous electrical loads 32. EDGIE implements empirically validated, multi-physics modeling at hourly time resolution and county-level spatial resolution. We tune building, appliance, and vehicle parameters to county-specific data. We find that almost all United States counties could require grid reinforcement in a 100% electrification scenario. In colder climate zones, peak demand rises by 200% on average relative to today’s peak demand, and up to 800% in the coldest counties under the worst-case equipment choices. Distribution grid reinforcement to accommodate 100% electrification nationwide could cost \$320 to \$720 billion, or about \$2,500 to \$5,800 per household. However, demand-side management strategies, such as improving equipment efficiencies and coordinating device operation, could reduce nationwide grid reinforcement costs by up to 77%.
RESULTS
@@
We assess the physical impacts of home and vehicle electrification on distribution grids through bottom-up modeling. In each county, we simulate all the residential appliances and personal vehicles from 1,000 representative households, then rescale the aggregate load profile to reflect the true number of households. Simulated appliance and vehicle properties, housing sizes and types (detached single-family vs. attached multifamily), and levels of insulation and air sealing all vary both across and within counties. We simulate peak heating and cooling weeks under two scenarios: (1) A business-as-usual scenario that reflects today’s building stock, appliances, and vehicles, and (2) an all-electric scenario that reflects a possible future with complete electrification of personal vehicles and residential space and water heating. In the all-electric scenario, each home has a resistance water heater, an air-to-air heat pump for space heating and cooling, and backup resistance heat. We compute design heating and cooling loads for each home under EnergyStar heating and cooling design conditions 33. Each heat pump is sized to meet the larger of its design heating load and design cooling load, capped at the largest residential air-to-air heat pump size currently available. In the two northernmost climate zones, we simulate adoption of heat pumps that meet the United States Department of Energy’s Cold Climate Heat Pump Challenge specifications 34. We fit physical and behavioral parameters associated with buildings, thermal equipment, and vehicles to county- and climate-specific data, using real data wherever possible. The Experimental Procedures section and the Supplementary Information document provide more details on models and data sources.
Figure 3: Today (top left), per-household demand peaks are typically higher in hotter areas and in cold rural areas with significant amounts of electric heating. In an all-electric future (top right), per-household peaks increase and shift to colder areas. Per-household grid reinforcement requirements (bottom) are highest in the Rocky Mountains, the upper Midwest, and the Northeast.
The modeling approach outlined above enables analysis of the sizes and causes of likely grid impacts from electrification. Fig. 2 shows the aggregate power used by 1,000 households over a peak cooling week in Miami, Florida (left plot; a hot, humid climate) and a peak heating week in Minneapolis, Minnesota (right; cold). In these plots, the shaded areas from bottom to top represent miscellaneous electrical loads (blue; lights, computers, dishwashers, etc.), water heaters (orange), electric vehicles (yellow), and air-source heat pumps with backup resistance heat (purple). Dashed red lines show today’s annual demand peaks. In Miami, electrification roughly doubles peak demand relative to today’s peak. Electric vehicle adoption increases electricity demand, but the new air-to-air heat pumps – which can both heat and cool – offset some of this increase by improving air conditioning efficiency. In Minneapolis, by contrast, peak demand increases by about a factor of five, driven mainly by space heating in the coldest weather, when heat demand rises and heating efficiencies fall.
@@
The lower heat map in Fig. 3 shows the estimated grid reinforcement capacity, normalized by the number of households, required to accommodate 100% electrification in each county. We estimate the grid reinforcement capacity requirement as the difference between the all-electric and business-as-usual grid capacities. As expected, northern regions require more grid reinforcement due to higher space heating demand. Grid reinforcement requirements reach a maximum of 24 kW per household in the northern Midwest. Many southern areas with mild winters require little or no grid reinforcement. In some cases, the lower heat map in Fig. 3 shows significant variation in grid reinforcement requirements between adjacent counties. This variation can be explained by differences in housing types and sizes, levels of insulation and air sealing, driving patterns, weather extremes, or other location-specific data. These geographic differences suggest a need for tailored infrastructure planning, as approaches that work in one county may not necessarily work in neighboring counties.
@@
To estimate the economic costs of reinforcing distribution grids to accommodate electrification, we multiply the aggregate grid reinforcement requirement for each county (in units of kW) by a \$/kW price. Informed by several recent studies, we choose this price to reflect a typical distribution grid upgrade. In reality, grid reinforcement prices vary significantly across the United States due to variation in grid topology and capacity, prevalence of overhead vs. underground power lines, costs of labor and equipment, and other factors. Due to a lack of reliable location-specific data on grid reinforcement prices, however, we use the same price distribution for each county.
@@
We draw on four studies 24, 35, 36, 37 to determine a plausible range of prices for typical distribution grid reinforcement, which may include installing power lines, transformers, switchgear, capacity banks, or other infrastructure. Based on Pacific Gas & Electric data, Elmallah et al. 24 estimate prices of 860 to 36,500 \$/kW, with the highest prices for the smallest projects. (We inflation-adjusted all price data to 2024 dollars.) Larson et al. 35 use Energy Information Administration data 38 to derive state-wise prices and a national average of 1,664 \$/kW, including a variety of administrative costs. Fares et al. 36 estimate an average price of 69 \$/kW based on data from the Federal Energy Regulatory Commission’s Form 1, where utilities can report their capital, operation, and maintenance costs related to grid reinforcement. In estimating this price, Fares et al. delineate distribution grid reinforcements from generation, transmission, and administrative costs. Rauschkolb et al. 37 use the same Federal Energy Regulatory Commission dataset to develop a linear model that accounts for both one-time capital costs of expanding grid capacity, and for increases in the annual costs of operating and maintaining a larger grid.
Here, we modify the approach of Rauschkolb et al. to estimate a representative price of distribution grid reinforcement. We assume that grid capacities increase linearly from the business-as-usual capacities in 2025 to the all-electric capacities in 2050. We map future costs into net present costs using discounted cash flow analysis. Similar to Rauschkolb et al.’s approach 37, our modified approach accounts for both one-time capital costs of reinforcing infrastructure, and increased annual costs from operating and maintaining a larger grid. The Experimental Procedures section contains the details of our modifications, which yield a mean grid reinforcement price of 960 \$/kW and a 95% confidence interval of 587 to 1,331 \$/kW.
The upper row of heat maps in Fig. 4 shows the estimated total distribution grid reinforcement cost for each of the lower 48 United States under low (left), medium (middle), and high (right) price scenarios. We compute these estimates by multiplying the grid reinforcement requirement, aggregated over all counties in the state, by prices of 587, 960, and 1,331 \$/kW, corresponding to the 2.5th, 50th, and 97.5th price percentiles. Unlike the per-household grid reinforcement requirement estimates in Fig. 3, the aggregate cost estimates in Fig. 4 reflect population density as well as load growth intensity. States with both high populations and high per-household grid reinforcement requirements have the highest total costs. For example, the 7.6 million households in New York, a densely populated state with cold winters, see a total cost of about \$56 billion, or about \$7,400 per household. Per-household grid reinforcement costs, shown in the bottom row of heat maps in Fig. 4, reach a maximum of \$19,000 per household under the 97.5th price percentile in the cold, sparsely populated state of North Dakota.
Aggregated over the lower 48 United States, distribution grid reinforcement to accommodate 100% electrification of homes and private vehicles could cost \$320 to \$720 billion, or \$2,600 to \$5,800 per household. For context, residential energy bills cost Americans about \$230 billion (\$1,800 per household) per year 39. The Inflation Reduction Act of 2022 allocated about \$370 billion (\$2,900 per household) over ten years to spending on energy and climate 40. The United States annual military budget was about \$820 billion (\$6,500 per household) in 2023 41.
@@
Figure 5: Switching from resistance to heat-pump water heaters (top left) has little influence on grid reinforcement requirements. Decreasing thermal demand from buildings (top right) and switching from air- to ground-source heat pumps (bottom right) reduce grid reinforcement requirements significantly, particularly in cold climates. Coordinating device operation (bottom left) reduces grid reinforcement requirements in all climates.
The estimated distribution grid reinforcement requirements and costs in the previous sections come from an electrification scenario that extrapolates forward today’s housing stock, vehicle fleet, equipment configurations and efficiencies, and user behavior. Unmodified, these electrification choices could require \$320 to \$720 billion in distribution grid reinforcement costs nationwide. However, electrification with strategic demand-side management – referred to here as ‘smart electrification’ – could reduce these costs substantially.
To assess the cost reduction potential of smart electrification, we simulate four demand-side management strategies. First, the resistance water heaters simulated above could shift to heat-pump water heaters, which use about half the electricity per unit of heat output. Second, thermal demand from buildings could decrease due to better insulation and air sealing, due to shifts from detached single-family housing to less energy-intensive attached multifamily housing, or due to shifts to smaller housing. Third, software could reduce peak aggregate demand by coordinating the operation of water heaters, electric vehicles, and heat pumps. For example, heat pumps could preheat homes in anticipation of aggregate demand peaks, or electric vehicles could interleave their charging with heat pump operation. (We simulate device coordination through convex optimization with perfect information in one representative location per state, as discussed in the Supplementary Information document.) Fourth, the air-source heat pumps simulated above could shift to ground-source heat pumps, which use about one-third less electricity per unit of heat output and require less backup resistance heat.
@@
Fig. 6 shows nationwide grid reinforcement cost reductions from the four individual strategies and from combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$522 billion without demand-side management. Individually (top set of bars), the four strategies reduce costs by 4%, 13%, 36%, and 43%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication, rather than equipment upgrades – reduces costs more than switching to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of heat-pump water heaters (4% individually), less thermal demand from buildings (13%), and ground-source heat pumps (36%) reduces costs by 46%. The combination of all four demand-side management strategies reduces nationwide grid reinforcement costs by 77%, to a total of \$119 billion (\$950 per household).